OpenStudy (anonymous):

A rectangular piece of paper with length 31cm and width 14cm has two semicircles cut out of it, as shown below. Find the area of the paper that remains. Use the value 3.14 for π , and do not round your answer. Be sure to include the correct unit in your answer.

OpenStudy (anonymous):

@amoodarya

OpenStudy (anonymous):

Okay. So first step we have to find the area of the rectangular piece of paper. Formula for the area of a rectangle is A = l*h

OpenStudy (anonymous):

so the area is 544?

OpenStudy (anonymous):

Next step. Find the area of the semicircles. Because they make one big circle, we can just find the area of this circle. Area for a circle is: A = pi * r^2 The radius will be as kantalope said 8 because the radius is half of the diameter.

OpenStudy (anonymous):

Okay well before we go into details I know how to work the problem out I just am not sure of when to subtract or add

OpenStudy (anonymous):

We know the diameter because they are full half circles off the edge of the paper. That makes the diameter the same as the height of the paper

OpenStudy (anonymous):

which is 200.96

OpenStudy (anonymous):

Now you just subtract the area of the circle from the area of the rectangle: Area left = 544 - 200.96

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

okay, so the answer would be 343.04 cm^2?

OpenStudy (anonymous):

nope told me it was wrong

OpenStudy (anonymous):

Thats because he copied and pasted someone else's work lol http://openstudy.com/study#/updates/5164c953e4b0902dbc1cc94f

OpenStudy (anonymous):

Oh dang!! Lol do you know anything about these problems I had someone helping me and they were doing a good job and I was getting them right I just need help cause I cant get past it

OpenStudy (anonymous):

Im actually trying right now (:

OpenStudy (anonymous):

thank you :) I just don't understand here is a fresh one that I am working on right now if you want to start with that

OpenStudy (anonymous):

A training field is formed by joining a rectangle and two semicircles, as shown below. The rectangle is 91m long and 62m wide. Find the area of the training field. Use the value 3.14 for π , and do not round your answer. Be sure to include the correct unit in your answer.

OpenStudy (anonymous):

@hockeybeast10 @Electronia anyone else know how to help with this problem

OpenStudy (anonymous):

First you would have to multiply 91m * 2 = 182m

OpenStudy (anonymous):

yes I know that and then you have to divide 62 by 2

OpenStudy (anonymous):

any luck

OpenStudy (anonymous):

Remind me when where pie comes in, i completely forgot

OpenStudy (anonymous):

None :( I completely forgot how these problems were done :((((((

OpenStudy (anonymous):

oh ok well thanks anyways

OpenStudy (anonymous):

The equations of three lines are given below. Line 1: =y−−23x6 Line 2: =−6x4y−2 Line 3: =−2y+3x7 For each pair of lines, determine whether they are parallel, perpendicular, or neither.

OpenStudy (anonymous):

do you know anything about these

OpenStudy (anonymous):

@mooniepie

OpenStudy (anonymous):

any luck?

OpenStudy (anonymous):

line 1 nethier

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

What about line 2 and 3

OpenStudy (anonymous):

In numbers, if the one line's slope is m = 4/5, then the perpendicular line's slope will be m = –5/4. If the one line's slope is m = –2, then the perpendicular line's slope will be m = 1/2. To answer this question, I'll find the slopes. Since these two lines have identical slopes, then these lines are parallel.

OpenStudy (benlindquist):

2 and 3

OpenStudy (anonymous):

i do not know

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

well i need someones help its my last topic and i need to get it done

OpenStudy (anonymous):

@gabby1214 can you help